# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ 4.2 ChebTuck approximation for multi-particle potentials
In this section, we first recall the construction of the range-separated tensor format [4] for the total potential Eq. 4.1 in a $N$-particle system. Then we consider two practical examples, a bio-molecular system and a lattice-structured compound with vacancies, where we apply the ChebTuck approximation to the long-range part of the total potential.
The total potential of a $N$-particle system $P(x)$ in Eq. 4.1 is nothing but a weighted sum of shifted Newton kernels. It was proven in [25] that the tensor representation of the colocation tensor of the total potential can also be obtained by shifting the colocation tensor of the Newton kernel but with a more involved tweak. In fact, it can be obtained with help of shifting and windowing operator applied to the reference potential in Eq. 4.4, see [25],
${\cal W}_{\nu}={\cal W}_{\nu}^{(1)}\otimes{\cal W}_{\nu}^{(2)}\otimes{\cal W}_{\nu}^{(3)},\quad\nu=1,...,N,$
which applies to every particle located at $x_{\nu}$ in an $N$-particle system. Numerically it is obtained by construction of the single generating (reference) potential $\widetilde{\bf P}_{R}\in\mathbb{R}^{2n\times 2n\times 2n}$ in a twice larger computational box and shifting it according to $x_{1},x_{2},x_{3}$ coordinates of every particle with further restricting (windowing) it to the original computational domain. Then the total multi-particle potential in Eq. 4.1 is computed as a projected weighted sum of shifted and windowed potentials (see also [26] for more details),
${\bf P}_{0}=\sum_{\nu=1}^{N}{z_{\nu}}\,{\cal W}_{\nu}(\widetilde{\bf P}_{R})\in\mathbb{R}^{n\times n\times n}.$ (4.9)
It was proven in [4] that tensor representation of type (4.9) in case of a multi-particle system of general type, has almost irreducible large rank, $R_{N}=O(NR)$. To overcome this drawback, in [4] a new RS tensor format was introduced based on separation of the long- and short-range skeleton vectors in the generating tensor $\widetilde{\bf P}_{R}=\widetilde{\bf P}_{R_{s}}+\widetilde{\bf P}_{R_{l}}$ similar to the single Newton kernel version as in Section 4.1.3 which leads to separable representation for the total sum of potentials,
@@
Thanks to the RS tensor format, the CP-rank of the part of the sum, $\mathbf{P}_{l}$, containing only the long range components of the canonical vectors of the shifted kernels, scales as $R_{l}=O(R\log N)$, i.e., logarithmically in the number of particles in a molecular system [4]. Such rank compression of the long range part $\mathbf{P}_{l}$ is achieved by applying the RHOSVD method for the canonical to Tucker (C2T) transform, and the subsequent Tucker to canonical (T2C) decomposition [31, 28]. This yields the resulting canonical tensor $\mathbf{P}_{R_{l}}$, ${\mathbf{P}}_{l}\mapsto{\mathbf{P}}_{R_{l}}$ with the reduced rank $RN\mapsto R_{l}=O(R\log N)$.
One can aggregate the sum of short range tensors ${\bf P}_{s}$ (having local supports which do not overlap each other) into the so called “cumulated canonical tensors”, $\mathbf{P}_{s}$, parametrized only by the coordinates of particle centers and parameters of the localized reference tensor, $\widetilde{\mathbf{P}}_{R_{s}}$. Above considerations resulted in the following definition (see [4] for more details).
@@
(RS-canonical tensors [4]). Given a reference tensor ${\bf A}_{0}$ supported by a small box such that $\mbox{rank}({\bf A}_{0})\leq R_{0}$, the separation parameter $\gamma\in\mathbb{N}$, and a set of distinct points ${\bf x}_{\nu}\in\mathbb{R}^{d}$, $\nu=1,\cdots,N$, the RS-canonical tensor format specifies the class of $d$-tensors ${\bf A}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, which can be represented as a sum of a rank-${R}_{l}$ canonical tensor
${\bf A}_{R_{l}}={\sum}_{k=1}^{R_{L}}\xi_{k}{\bf a}_{k}^{(1)}\otimes\cdots\otimes{\bf a}_{k}^{(d)}\in\mathbb{R}^{n_{1}\times...\times n_{d}},$ (4.10)
and a cumulated canonical tensor
$\widehat{\bf A}_{S}={\sum}_{\nu=1}^{N}c_{\nu}{\bf A}_{\nu},$ (4.11)
generated by replication of the reference tensor ${\bf A}_{0}$ to the points ${\bf x}_{\nu}$. Then the RS canonical tensor is represented in the form
${\bf A}={\bf A}_{R_{l}}+\widehat{\bf A}_{S}={\sum}_{k=1}^{R_{L}}\xi_{k}{\bf a}_{k}^{(1)}\otimes\cdots\otimes{\bf a}_{k}^{(d)}+{\sum}_{\nu=1}^{N}c_{\nu}{\bf A}_{\nu},$ (4.12)
where $\mbox{diam}(\mbox{supp}{\bf A}_{\nu})\leq 2\gamma$ in the index size.
Figure 4.4: Left: Middle slice ${\bf A}_{R_{l}}(:,:,n/2)$ of the (compressed) long-range part of the RS tensor for the electrostatic potential of a biomolecule with $N=500$ particles computed on the 3D $n\times n\times n$ Cartesian grid with $n=1024$. with the canonical tensor rank $R_{l}=506$. Middle: the error of the canonical-to-Tucker transform, with $\varepsilon=10^{-6}$. Right: the total error of the canonical-to-canonical tensor transform with the $\varepsilon=10^{-5}$.
In Fig. 4.4 we visualize the middle slice for the long-range part of the RS tensor, see the left figure, representing the total electrostatic potential of a biomolecule with $N=500$ particles, i.e., ${\bf A}_{R_{l}}$ of the $N$-particle potential in CP tensor format, with the reduced canonical rank $R_{l}=506\sim R\log(N)$ from the original canonical rank $R_{N}=7000\sim NR$ computed by RHOSVD-based rank reduction method [31]. We also show the error of the canonical-to-Tucker transform and of the consequent Tucker-to-canonical transform, ultimately canonical-to-canonical transform, in the middle and right figures in Fig. 4.4, respectively. We observe a good approximation error recovering the accuracy of rank truncation in tensor transforms. In Fig. 4.5, we also visualize the several canonical vectors ${\bf a}_{k}^{(1)}$, $k=1,\cdots,10$ in the $x$-direction of the long-range part ${\bf A}_{R_{l}}$ of the same potential as in Fig. 4.4.
@@
Figure 4.5: Canonical vectors ${\bf a}_{k}^{(1)}$ of the long-range part in the total potential ${\bf A}_{R_{l}}\in\mathbb{R}^{n\times n\times n}$ with $N=500$ particles and grid size $n=256$ for $k=1,\cdots,8$.
4.2.1 A bio-molecule potential with $N=500$ particles
@@
In this section, we consider the ChebTuck approximation to the long range part of a bio-molecule with $N=500$ particles. In this case, the underlining function we want to approximate is given by a 3-dimensional canonical tensor ${\bf A}_{R_{l}}\in\mathbb{R}^{n\times n\times n}$, which can be pre-computed for any given grid size $n$ as described in Section 4.2. In this case, Algorithm 1 is not applicable but Algorithms 2, 3 and 4 can be used.
@@
We first consider the runtime and accuracy of Algorithms 2, 3 and 4 for different discretization grid sizes $n$. We set the Chebyshev degrees in each dimension to be $m_{1}=m_{2}=m_{3}=m=129$ and the Tucker-ALS (RHOSVD in case of Algorithm 4) truncation tolerance to be $10^{-6}$. We record the runtime and $\ell_{\infty}$ error at the middle slice for each case in Tables 4.4 and 4.5. Note that since the trivariate cubic spline in Algorithm 2 is cubic in each spatial direction and for each Chebyshev degree $m$ (see Proposition 3.11), we only compute the results for $n=256$. We observe that the accuracy of the three algorithms is similar. In particular, the Chebyshev interpolants produced by Algorithms 3 and 4 are mathematically equivalent before the Tucker compression, see also Remark 3.3. However, Algorithm 4 is the most efficient which leverages the CP tensor structure of the long-range part as well as the RHOSVD method for Tucker compression. This is consistent with the theoretical estimates in Proposition 3.11. The ChebTuck format also becomes more accurate with increasing grid size $n$ which provides more information for the underlying function to be approximated.
Next, we visualize the ChebTuck approximation error as well as the RHOSVD compression error for the potential with $n=256$ as in Fig. 4.6. In the left figure, we show the middle slice of the approximated potential $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$. Note that here we abuse the notation and use $\hat{f}_{\bf m}(:,:,t_{n/2})$ to denote the matrix containing function values of $\hat{f}_{\bf m}$ at the uniform grid points in the middle slice, i.e., $\hat{f}_{\bf m}(:,:,t_{n/2})=[\hat{f}_{\bf m}(t_{i_{1}},t_{i_{2}},t_{n/2})]_{i_{1}i_{2}}$. In the middle and right figures, we show the error during RHOSVD compressing and the total ChebTuck approximation error, i.e., $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$ and $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$, respectively. Note the difference between $\hat{f}_{\bf m}$ and $\tilde{f}_{\bf m}$ as described in Section 3.1. We observe good approximation accuracy of our ChebTuck format. Another notable observation is that the error during RHOSVD compression is much smaller than the prescribed tolerance $\varepsilon=10^{-7}$, which indicates that the singular values decay extremely fast after the first few significant ones. Similar figures are also obtained for larger grid sizes $n=512,1024,2048$ as shown in Fig. 4.7. Our ChebTuck format achieves even higher accuracy for larger grid sizes, consistent with the results in Table 4.5.
Algorithm 2 Algorithm 3 Algorithm 4
$n=256$ 13.52 0.57 0.14
$n=512$ 0.48 0.20
$n=1024$ 0.50 0.28
$n=2048$ 0.55 0.25
@@
Table 4.4: Runtime in seconds for the ChebTuck approximation of the potential of a bio-molecule with $N=500$ particles in $\mathbb{R}^{3}$ for different grid sizes $n$.
Algorithm 2 Algorithm 3 Algorithm 4
$n=256$ $8.59\cdot 10^{-7}$ $8.59\cdot 10^{-7}$ $8.59\cdot 10^{-7}$
$n=512$ $6.89\cdot 10^{-8}$ $6.89\cdot 10^{-8}$
$n=1024$ $4.21\cdot 10^{-9}$ $4.21\cdot 10^{-9}$
$n=2048$ $6.37\cdot 10^{-11}$ $6.37\cdot 10^{-11}$
Table 4.5: $\ell_{\infty}$ error for the ChebTuck approximation in middle slice of the potential, i.e., $\max_{i_{1},i_{2}}|{\bf A}_{R_{l}}(i_{i},i_{2},n/2)-\hat{f}_{{\bf m}}(t_{i_{1}},t_{i_{2}},t_{n/2})|$ for different grid sizes $n$.
@@
We also remark here that the choice of the univariate interpolations $q_{k}^{(\ell)}(x)$ in Eq. 3.12 is crucial for the accuracy of the ChebTuck approximation since the convergence rate of the Chebyshev interpolants $g_{k}^{(\ell)}(x)$ depends on the smoothness of the univariate interpolations $q_{k}^{(\ell)}(x)$. For example, in Fig. 4.8, we replace the spline interpolation in Eq. 3.12 by piecewise linear (middle) and nearest neighbor (right) interpolation and observe that the ChebTuck format accuracy deteriorates significantly even for a large grid size $n=2048$.
@@
Figure 4.6: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck format of a bio-molecule with $N=500$ and $n=256$. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
@@
Figure 4.7: The degree $m=129$ ChebTuck approximation total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$ of the same bio-molecule for grid sizes $n=512,1024,2048$ from left to right.
@@
Figure 4.8: The total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck approximation of the same bio-molecule for grid size $n=2048$ with different univariate interpolations in Eq. 3.12. Left: cubic spline, middle: piecewise linear, right: nearest neighbor.

### v2

@@ 4.2 ChebTuck approximation for multi-particle potentials
In this section, we first recall the construction of the range-separated tensor format [6] for the total potential Eq. 33 in an $N$-particle system. Then we consider two practical examples, a bio-molecular system and a lattice-structured compound with vacancies, where we apply the ChebTuck approximation to the long-range part of the total potential.
The total potential of an $N$-particle system $P(x)$ in Eq. 33 is nothing but a weighted sum of shifted Newton kernels. It was proven in [33] that the tensor representation of the collocation tensor of the total potential can also be obtained by shifting the collocation tensor of the Newton kernel, albeit through a more elaborate construction. In fact, it can be obtained through the application of shifting and windowing operators to the reference potential in Eq. 35, see [33], ${\cal W}_{\nu}={\cal W}_{\nu}^{(1)}\otimes{\cal W}_{\nu}^{(2)}\otimes{\cal W}_{\nu}^{(3)}$, $\nu=1,\cdots,N$, which applies to every particle located at $x_{\nu}$ in an $N$-particle system. Numerically it is obtained by construction of the single generating (reference) potential $\widetilde{\bf P}_{R}\in\mathbb{R}^{2n\times 2n\times 2n}$ in a twice larger computational box and shifting it according to the $x_{1},x_{2},x_{3}$ coordinates of every particle with further restricting (windowing) it to the original computational domain. Then the total multi-particle potential in Eq. 33 is computed as a projected weighted sum of shifted and windowed potentials ${\bf P}_{0}=\sum_{\nu=1}^{N}{z_{\nu}}\,{\cal W}_{\nu}(\widetilde{\bf P}_{R})\in\mathbb{R}^{n\times n\times n}$, see also [34] for more details.
It was proven in [6] that a tensor representation of this type in case of a multi-particle system of general type has almost irreducible large rank, $R_{N}=O(NR)$. To overcome this drawback, in [6] a new RS tensor format was introduced based on separation of the long- and short-range skeleton vectors in the generating tensor $\widetilde{\bf P}_{R}=\widetilde{\bf P}_{R_{s}}+\widetilde{\bf P}_{R_{l}}$ similar to the single Newton kernel version as in Section 4.1.3 which leads to a separable representation for the total sum of potentials,
@@
Thanks to the RS tensor format, the CP-rank of the part of the sum, $\mathbf{P}_{l}$, containing only the long-range components of the canonical vectors of the shifted kernels, scales as $R_{l}\in O(R\log N)$, i.e., logarithmically in the number of particles in a molecular system [6]. Such rank compression of the long-range part $\mathbf{P}_{l}$ is achieved by applying the RHOSVD method for the canonical to Tucker (C2T) transform, and the subsequent Tucker to canonical (T2C) decomposition [39, 38]. This yields the resulting canonical tensor $\mathbf{P}_{R_{l}}$, ${\mathbf{P}}_{l}\mapsto{\mathbf{P}}_{R_{l}}$ with the reduced rank $RN\mapsto R_{l}\in O(R\log N)$.
One can aggregate the sum of short-range tensors ${\bf P}_{s}$ (having local supports which do not overlap each other) into the so called “cumulated canonical tensors”, $\mathbf{P}_{s}$, parametrized only by the coordinates of particle centers and parameters of the localized reference tensor, $\widetilde{\mathbf{P}}_{R_{s}}$. Above considerations resulted in the following definition (see [6] for more details).
@@
(RS-canonical tensors [6]). Given a reference tensor ${\bf A}_{0}$ supported by a small box such that $\mbox{rank}({\bf A}_{0})\leq R_{0}$, the separation parameter $\gamma\in\mathbb{N}$, and a set of distinct points ${\bf x}_{\nu}\in\mathbb{R}^{d}$, $\nu=1,\ldots,N$, the RS-canonical tensor format specifies the class of $d$-tensors ${\bf A}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$, which can be represented as a sum of a rank-${R}_{l}$ canonical tensor ${\bf A}_{R_{l}}$ and a cumulated canonical tensor $\widehat{\bf A}_{S}$,
(39) ${\bf A}_{R_{l}}={\sum}_{k=1}^{R_{L}}\xi_{k}{\bf a}_{k}^{(1)}\otimes\cdots\otimes{\bf a}_{k}^{(d)},\ \widehat{\bf A}_{S}={\sum}_{\nu=1}^{N}c_{\nu}{\bf A}_{\nu}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}},$
where $\widehat{\bf A}_{S}$ is generated by replication of the reference tensor ${\bf A}_{0}$ to the points ${\bf x}_{\nu}$. Then the RS canonical tensor is represented in the form ${\bf A}={\bf A}_{R_{l}}+\widehat{\bf A}_{S}$, where $\mbox{diam}(\mbox{supp}{\bf A}_{\nu})\leq 2\gamma$ in the index size.
In what follows, we consider biomolecular systems corresponding to truncated fragments of the protein Fasciculin 1 with different numbers of atoms $N$. Fasciculin 1 is an anti-acetylcholinesterase toxin from the green mamba snake venom comprising 1,228 atoms [5, 41]. Throughout this section and Sections 4.2.1 and 4.2.2, the term biomolecule refers to this specific system.
In Fig. 4, we visualize the middle slice for the long-range part of the RS tensor, see the left figure, representing the total electrostatic potential of the biomolecule with $N=500$ particles, i.e., ${\bf A}_{R_{l}}$ of the $N$-particle potential in CP tensor format, with the reduced canonical rank $R_{l}=506\sim R\log(N)$ as compared to the original canonical rank $R_{N}=7000\sim NR$ computed by the RHOSVD-based rank reduction method [39]. We also show the error of the canonical-to-Tucker transform and of the consequent Tucker-to-canonical transform, ultimately canonical-to-canonical transform, in the middle and right figures in Fig. 4, respectively. We observe a good approximation error recovering the accuracy of rank truncation in tensor transforms. In Figure 1 right, we also visualize the several canonical vectors ${\bf a}_{k}^{(1)}$, $k=1,\ldots,10$, in the $x$-direction of the long-range part ${\bf A}_{R_{l}}$ of the same potential as in Figure 4.
@@
Figure 4: Left: Middle slice ${\bf A}_{R_{l}}(:,:,n/2)$ of the (compressed) long-range part of the RS tensor for the electrostatic potential of the biomolecule with $N=500$ particles computed on the 3D $n\times n\times n$ Cartesian grid with $n=1024$ and rank $R_{l}=506$. Middle: the error of the canonical-to-Tucker transform, with $\varepsilon=10^{-6}$. Right: the total error of the canonical-to-canonical tensor transform with the $\varepsilon=10^{-5}$.
4.2.1 The biomolecule potential with $N=500$ particles
@@
In this section, we consider the ChebTuck approximation to the long range part of the biomolecule with $N=500$ particles. In this case, the underlining function we want to approximate is given by a 3-dimensional canonical tensor ${\bf A}_{R_{l}}\in\mathbb{R}^{n\times n\times n}$, which can be pre-computed for any given grid size $n$ as described in Section 4.2. In this case, Algorithm 1 is not applicable but Algorithms 2 and 3 can be used.
@@
We first consider the runtime and accuracy of Algorithms 2 and 3 for different discretization grid sizes $n$. We set the Chebyshev degrees in each dimension to be $m_{1}=m_{2}=m_{3}=m=129$ and the Tucker-ALS (RHOSVD in case of Algorithm 3) truncation tolerance to be $10^{-6}$. We record the runtime and $\ell_{\infty}$ error at the middle slice for each case in Table 3. Note that since the trivariate cubic spline in Algorithm 2 is cubic in grid size $n$ and for each Chebyshev degree $m$ (see Proposition 3.17), we only compute the results for $n=256$. We observe that the accuracy of the algorithms is similar while Algorithm 3 is significantly more efficient as it leverages the CP tensor structure of the long-range part as well as the RHOSVD method for Tucker compression. This is consistent with the theoretical estimates in Proposition 3.17. The ChebTuck format also becomes more accurate with increasing grid size $n$ which provides more information for the underlying function to be approximated.
Algorithm 2 Algorithm 3
$\ell_{\infty}$ error runtime $\ell_{\infty}$ error runtime
$n=256$ $8.59\cdot 10^{-7}$ 9.44 $8.59\cdot 10^{-7}$ 0.07
$n=512$ $6.89\cdot 10^{-8}$ 0.09
$n=1024$ $4.21\cdot 10^{-9}$ 0.13
$n=2048$ $6.37\cdot 10^{-11}$ 0.18
@@
Table 3: Runtime in seconds and $\ell_{\infty}$ error for the ChebTuck approximation of the potential of the biomolecule with $N=500$ particles in $\mathbb{R}^{3}$ for different grid sizes $n$.
Next, we visualize the ChebTuck approximation error as well as the RHOSVD compression error for the potential with $n=256$ as in Fig. 5. In the left figure, we show the middle slice of the approximated potential $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$. Note that here we abuse the notation and use $\hat{f}_{\bf m}(:,:,t_{n/2})$ to denote the matrix containing function values of $\hat{f}_{\bf m}$ at the uniform grid points in the middle slice, i.e., $\hat{f}_{\bf m}(:,:,t_{n/2})=[\hat{f}_{\bf m}(t_{i_{1}},t_{i_{2}},t_{n/2})]_{i_{1}i_{2}}$. In the middle and right figures, we show the error during RHOSVD compressing and the total ChebTuck approximation error, i.e., $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$ and $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$, respectively. Note the difference between $\hat{f}_{\bf m}$ and $\tilde{f}_{\bf m}$ as described in Section 3.1. We observe good approximation accuracy of our ChebTuck format. Another notable observation is that the error during RHOSVD compression is much smaller than the prescribed tolerance $\varepsilon=10^{-7}$, which indicates that the singular values decay extremely fast after the first few significant ones. Similar figures are also obtained for larger grid sizes $n=512,1024,2048$ as shown in Fig. 6. Our ChebTuck format achieves even higher accuracy for larger grid sizes, consistent with the results in Table 3.
@@
We also remark here that the choice of the univariate interpolations $q_{k}^{(\ell)}(x)$ in Eq. 19 is crucial for the accuracy of the ChebTuck approximation since the convergence rate of the Chebyshev interpolants $g_{k}^{(\ell)}(x)$ depends on the smoothness of the univariate interpolations $q_{k}^{(\ell)}(x)$, as discussed in Lemma 3.8. For example, in Fig. 7, we replace the spline interpolation in Eq. 19 by piecewise linear (middle) and nearest neighbor (right) interpolation and observe that the ChebTuck format accuracy deteriorates significantly even for a large grid size $n=2048$.
@@
Figure 5: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck format of the biomolecule with $N=500$ and $n=256$. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
@@
Figure 6: The degree $m=129$ ChebTuck approximation total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$ of the biomolecule for grid sizes $n=512,1024,2048$ from left to right.
@@
Figure 7: The total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck approximation of the biomolecule for grid size $n=2048$ with different univariate interpolations in Eq. 19. Left: cubic spline, middle: piecewise linear, right: nearest neighbor.